OpenStudy (anonymous):

What am I missing from my answers below this question? Use mathematical induction to prove the statement is true for every positive integer n. 8+16+24+...+8n=4n(n+1) ANSWER: 8n=4n(n+1) 8(1)=4(1)(1+1) 8=8 (Proved correct) ANSWER: 4k(k+1)=8 4(1) (1+1) = 8 4 (2) = 8 8 = 8 (Proved correct.

OpenStudy (xapproachesinfinity):

Mathematical induction requires you to check if P(0) is true then You assume P(n) is true, you have to show that if P(n) is true then P(n+1) is also true therefore your statement as is true

OpenStudy (anonymous):

So my answer then is correct.

OpenStudy (xapproachesinfinity):

well not really! you have proven that for n=1 your statement is valid but where is the assumption that P(n) is true then prove that p(n+1) is true?

OpenStudy (xapproachesinfinity):

when you do P(n) and P(n+1) you don't check for specific examples you need to prove for general case

OpenStudy (anonymous):

Oh geez. Help then please?\

OpenStudy (xapproachesinfinity):

Okay! Assume P(n) is true that is saying that 8+16+24+....+8n=4n(n+1) is true

OpenStudy (xapproachesinfinity):

now we check that if P(n+1) is true or not so P(n+1): 8+16+24+....+8(n+1)=4(n+1)(n+2)

OpenStudy (xapproachesinfinity):

So far good , yes?

OpenStudy (anonymous):

Hold on one moment, so I can review what you have shown me.

OpenStudy (xapproachesinfinity):

eh i have to go to class so be quick hehhe

OpenStudy (anonymous):

we added (n+2) to the equation

OpenStudy (xapproachesinfinity):

No we just plug in n+1 in place of n so we can get n+1 expression and we need to show that indeed is equal and true in P(n+1) i put = sign but we are still asking is this true? you got me the point is to show that left hand side is indeed equal right hand side

OpenStudy (xapproachesinfinity):

we going to use our assumption that P(n) is true that we assumed!

OpenStudy (anonymous):

Got it. And that I believe would be the only equation I am missing . I knew it. Thanks.

OpenStudy (xapproachesinfinity):

okay! Actually you missed to the assumption too cause if you didn't do this step you are not going to solve this

OpenStudy (xapproachesinfinity):

And you are welcome^_^

OpenStudy (anonymous):

I got it. Thanks and have a good class.

OpenStudy (xapproachesinfinity):

Welcome mate! good luck with this to better get the hang of it practice more example with different settings

OpenStudy (anonymous):

Am doing that, but after a few hours the brain freezes and you begin doubting work you've done. Have to learn to take breaks in between! Thanks again.